OpenStudy (anonymous):

If I = prt, the value of I when p = 1,000, r = 0.05, and t = 3 is a0. Help, I don't get this question at all.

jimthompson5910 (jim_thompson5910):

I = p*r*t I = 1000*0.05*3 ... plug in p = 1,000, r = 0.05, and t = 3 I = 50*3 I = 150

jimthompson5910 (jim_thompson5910):

so you plug in the given values, then evaluate

jimthompson5910 (jim_thompson5910):

in real world terms this means that if you deposit $1,000 at 5% simple interest for 3 years, you earn $150 in interest alone.

OpenStudy (anonymous):

I see what you did now, thank you! I understand it now.

jimthompson5910 (jim_thompson5910):

that's great